# WebCoderBench: Benchmarking Web Application Generation with Comprehensive and Interpretable Evaluation Metrics

### v1

@@ Abstract
Chenxu Liu1thanks: The email address of the first author is: chenxuliu@stu.pku.edu.cn, Yingjie Fu1, Wei Yang3, Ying Zhang2, Tao Xie1thanks: Tao Xie is with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education; School of Computer Science, Peking University, Beijing, China; Fudan University Institute of Systems for Advanced Computing, Shanghai, China; Shanghai Institute of Systems for Open Computing, Shanghai, China. 1Key Lab of HCST (PKU), MOE; SCS, Peking University, China 2Key Lab of HCST (PKU) & NERCSE, MOE; Peking University, China 3University of Texas at Dallas, USA Correspondence: taoxie@pku.edu.cn
@@ 1 Introduction
Existing benchmarks for evaluating LLMs on web app generation fall into three main categories, each facing limitations in addressing the aforementioned challenges. First, synthetic benchmarks webgenbench; fullfront; frontendbench; webbench; artifactsbench construct datasets using LLM-generated or expert-written natural language requirements, failing to capture the diversity and authenticity of real-world user expressions. Second, reference-based benchmarks pix2code; design2code; sketch2code; vision2ui; web2code; webcode2m require LLMs to reproduce web apps from provided screenshots or sketches, thus lacking the generality needed to evaluate open-ended user requirements. Third, arena-style benchmarks WebDevArena; designbench collect real-world user requirements and rank LLMs through large-scale human voting in blind evaluations arenaranking. While these results align with human preferences, such benchmarks depend heavily on manual annotations and fail to deliver fine-grained, quantitative evaluations for a deep analysis.
@@ 2.1 LLMs for Web App Generation
The LLMs for web app generation mainly fall into three categories. The first category is foundation models, including general-purpose models (e.g., GPT5 GPT, Claude Claude, Gemini Gemini, DeepSeek DeepSeek, Qwen Qwen3, Doubao Doubao) and code-specialized models (e.g., Deepseek-Coder deepseekcoder, Qwen-Coder Qwen2.5coder, StarCoder starcoder). These models process multi-language code synthesis, long-context reasoning, and iterative repair via execution feedback. The second category is IDE/CLI-centric coding agents (e.g, Cursor Cursor, Trae TRAE, Windsurf Windsurf, and Claude Code ClaudeCode). These agents support manipulating local projects through project bootstrapping, dependency resolution, multi-file editing, and command-line tool control. The third category is chat-based coding agents (e.g., Manus Manus, MiniMax agent Minimax-agent, Doubao Coding DoubaoCoding, and Lovable Lovable).
@@ 2.2 Benchmarks for Web App Generation
WebGen-Bench webgenbench Text 101 LLM + human experts Yes LLM as a judge + test case execution Test cases
FullFront fullfront Text + Image 1,800 LLM Yes Code & visual similarity + LLM as a judge Code + Images
FrontendBench frontendbench Text 148 LLM Yes Test case execution Test cases
Web-Bench webbench Text 100 Human experts Yes Test case execution Test cases
WebDev Arena WebDevArena Text 10,501 Real user requirements No Pairwise manual labeling -
Design Arena DesignArena Text - Real user requirements No Pairwise manual labeling -
ArtifactsBench artifactsbench Text 1,825 LLM + human experts Yes LLM as a judge Checklist
@@
Web app generation has long been an active research topic. Beginning with pix2code pix2code, a large body of benchmarks and datasets websight; design2code; sketch2code; vision2ui; web2code; webcode2m focus on generating web apps from screenshots or sketches. These approaches typically conduct evaluation by comparing the generated web app code or rendered screenshots against the corresponding ground-truth code or original screenshots.
@@ 3.1 Data Collection
Next, we let each requirement be reviewed by three human experts to remove requirements that are incomprehensible, missing supplementary materials (e.g., images), or do not apply to native HTML scenarios. We further conduct text-level and semantic-level de-duplication using MinHash MinHash and MiniLM MiniLM semantic embedding, following ArtifactsBench artifactsbench. Our filtered dataset retains 1,572 requirements.
@@ 3.2 Dataset Statistics
We perform detailed and multi-dimensional classifications of each requirement to enable an in-depth analysis of the characteristics and distribution of our dataset. The classification results are shown in Figure 2, with all tags manually labeled by human experts assisted by an LLM (Gemini-2.5-pro).
@@
The artifact complexity dimension spans from highly simple to highly complex outputs, with most samples falling into medium (658) and simple (433) levels, showing balanced difficulty for model evaluation. In terms of input modality, our dataset mainly includes text-only requirements (1,413), complemented by text with images (123) and text with URLs (36), allowing evaluation of both uni-modal and multi-modal understanding abilities. It is worth noting that, unlike the task of generating code from screenshots pix2code; webcode2m, our user requirements can include images and URLs that are intended to serve as page content rather than as reference designs.
@@ 4 Evaluation Metrics
To comprehensively assess the quality of web app artifacts generated by LLMs, we design a set of evaluation metrics from multiple perspectives. Inspired by practices in the text-to-image domain text2imagesurvey, our evaluation considers two major aspects: general quality and alignment quality. These aspects are further divided into nine perspectives encompassing 24 detailed evaluation metrics, jointly providing a comprehensive evaluation. The metrics are defined based on public standards (e.g., Google Lighthouse Lighthouse, Lint HTMLHint; StyleLint; ESLint, W3C design principles W3C) and internal guidelines from our industrial partner (e.g., visual design and copywriting standards). We list all metrics in Table 3, with detailed information provided in Appendix A. Each metric produces a quantitative score ranging from 0 to 100, where a higher score indicates better quality.
@@ 4.3 Weight Assignment
After defining evaluation metrics, a crucial process is to combine the outcome scores in a meaningful way to generate an overall score, reflecting human preference. Our process of generating the overall score is shown in Figure 3. Existing studies webuibench; artifactsbench typically assign weights for each metric either uniformly or heuristically. However, real-world users do not treat all metrics equally and can prioritize certain perspectives (e.g., emphasizing code quality while paying less attention to maintainability). Consequently, heuristic weighting neither aligns with user priorities nor ensures fairness.
@@
To obtain preference-aligned weights for each metric, we conduct an internal survey within the company of our industrial partner. Specifically, we ask each participant to rank the 24 metrics according to their perceived importance. To reduce the cognitive burden on participants, each of them is first asked to rank the nine perspectives listed in Table 3, and then to rank the metrics within each perspective. During a three-day survey period, our questionnaire receives 1,076 views and 899 responses, corresponding to a response rate of 83.55%. We further filter the responses by completion time, retaining only those that take more than two minutes to complete. As a result, we obtain 141 valid responses. According to a recent study surveysample, a sample size of more than 100 responses can be considered as sufficient for ranking and regression analyses. Among the valid responses, there are 21 data scientists, 19 product managers, 22 legal personnel, 23 front-end/back-end developers, 2 designers, 39 operations staff, and 15 quality assurance personnel, according to their user personas.
We further apply the Borda Count borda algorithm to extract weights from responses. This algorithm assigns a score to each item (perspective or metric) based on its position in each participant’s ranking. The weight of each item is then obtained by summing its scores across all participants and normalizing by the total score over all items. Each metric’s weight is calculated by multiplying its own weight by the weight of the perspective it belongs to. The weights of all metrics sum to 1. The resulting weights are shown in Figure 4.
The overall score of each model is calculated by summing the z-scores zscore of all metrics, averaged over the 1,572 samples in the dataset, and weighted by the corresponding metric weights. We adopt z-scores to ensure that the scores of different metrics are on a comparable scale with consistent discriminability, thereby emphasizing the effect of the weights. The z-scores are calculated by $z_{i,j}=\frac{x_{i,j}-\mu_{j}}{\sigma_{j}}$, where $x_{i,j}$ denotes the raw score of the $i$-th sample on the $j$-th metric, and $\mu_{j}$ and $\sigma_{j}$ denote the mean and standard deviation of this metric over all samples, respectively.
@@ 5.1 Settings
We evaluate 12 representative LLMs and 2 LLM-based agents on WebCoderBench. The selected models span multiple families (e.g., DeepSeek DeepSeek, Qwen Qwen3, Gemini Gemini, GLM GLM, GPT GPT), multiple versions (e.g., base, instruct, thinking, coder), multiple modalities (e.g., text-only, multi-modal), and multiple capability types (e.g., standard Minimax-m1; Minimax-m2, agentic Manus; Minimax-agent).We invoke each LLM using its standard APIs with the recommended parameters and identical system prompts. For LLM-based agents, we manually enter prompts and requirements through their web interfaces to obtain the generated artifacts, and further constrain their outputs to native HTML by appending specific control prompts. Due to labor constraints, we collect artifacts for 165 requirements from these agents, corresponding to approximately one-tenth of the full dataset. To enable uni-modal LLMs (e.g., DeepSeek) to handle multi-modal requirements, we use Gemini-2.5-pro to generate textual descriptions for each image and provide these descriptions as input to the uni-modal LLMs. We let each model generate a single artifact for each requirement.
@@ 5.2 RQ1: Main Results
Compared to coder models (e.g., Qwen3-Coder-Plus and GPT-5-Codex), generalist models (e.g., Qwen3-Instruct and GPT-5-High) present better effectiveness, showing the task of web app generation requires not only coding ability, but also general understanding of requirements and real-world knowledge. This finding also aligns to ArtifactsBench artifactsbench.
@@ 6 Conclusion
Finally, metric correctness and stability can influence the evaluation. For rule‑based metrics, we manually inspect 50 scoring instances per metric and conduct additional code reviews. For LLM‑as‑a‑judge metrics, we run each metric three times on 50 requirements and use the Mann–Whitney U test mann1947test to verify that score variance across runs is substantially lower than variance across models. We follow prior work fullfront for prompt design and use different LLMs for different metrics to mitigate preference leakage and bias LLMjudgenotreliable; LLMjudgepreference; LLMjudgeuncertainty.
@@ F.5 Prompts for Evaluation Metrics
Prompts for the evaluation metric of Content Alignment: **Role-play:** You are a senior requirements analyst, project manager, and a user of web front-end applications. You have just received a web front-end application developed based on given requirements. It is still in the early stages of development. Your task is to determine whether the developer has correctly understood the user’s requirements and to judge at a high level whether the given web page code meets the user-specified requirement points, without worrying about the correctness of the specific implementation. **Task Goal:** Analyze the web front-end application code and determine if it meets the given requirement points. **Input:** **Web page code** and **requirement points** information **Evaluation Criteria:** Please analyze from the user’s perspective whether the code implements the specified requirement points based on the following dimensions: **Content Semantics:** Does the content on the page conform to the semantics specified in the user requirements? **Specified Data:** Does the content on the page include the specific data that the user specified to include? (e.g., image links, data source links, text copy, video/audio links, etc.) **Output:** Please output your judgment and analysis in JSON format, with the following structure: [{ "content_requirement": "<Requirement Point>", "code_snippet": "<Web Front-end Application Code> (The maximum output length is 1000 characters.)", "is_implemented": <true/false>, "implementation_analysis": "<A concise analysis of how the code implements the requirement point>", "confidence_score": <A confidence score from 0-1, indicating your level of certainty, with 1 being the highest> }] **Example:** **Input:** 1. **Code Snippet:** <div>Welcome</div> <img src="logo.png"> <a href="/home"></a> 2. **Requirement Points:** ["Display a welcome text.", "Display a company logo, using logo.png.", "Provide a link to return to the homepage."] **Expected Output (JSON format):** Note: The final output should only contain the JSON content, do not wrap it in Markdown’s JSON format. [{ "content_requirement": "Display a welcome text.", "code_snippet": "<div>Welcome</div>", "is_implemented": true, "implementation_analysis": "The code uses a ‘<div>’tag to contain the text Ẅelcome,̈ thus implementing the requirement to display a welcome text.", "confidence_score": 0.8 }, { "content_requirement": "Display a company logo.", "code_snippet": "<img src=l̈ogo.png>̈", "is_implemented": true, "implementation_analysis": "The code uses an ‘<img>’ tag to attempt to display the user-specified image named l̈ogo.png.̈", "confidence_score": 1.0 }, { "content_requirement": "Provide a link to return to the homepage.", "code_snippet": "<a href=/̈home>̈</a>", "is_implemented": false, "implementation_analysis": "The code uses an ‘<a>’ tag to create a link pointing to /̈home,̈ but there is no text on the link, which fails to indicate to the user that it is a link to return to the homepage.", "confidence_score": 0.9 }]

### v2

@@ Abstract
Chenxu Liu1thanks: Equal contribution.thanks: The email address of the first author is: chenxuliu@stu.pku.edu.cn, Yingjie Fu [^1] footnotemark: 1, Wei Yang3, Ying Zhang2, Tao Xie1thanks: Tao Xie is with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education; School of Computer Science, Peking University, Beijing, China; Fudan University Institute of Systems for Advanced Computing, Shanghai, China; Shanghai Institute of Systems for Open Computing, Shanghai, China. 1Key Lab of HCST (PKU), MOE; SCS, Peking University, China 2Key Lab of HCST (PKU) & NERCSE, MOE; Peking University, China 3University of Texas at Dallas, USA Correspondence: taoxie@pku.edu.cn
@@ 1 Introduction
Existing benchmarks for evaluating LLMs on web app generation fall into three main categories, each facing limitations in addressing the aforementioned challenges. First, synthetic benchmarks Lu et al. (2025); Sun et al. (2025); Zhu et al. (2025); Xu et al. (2025); Zhang et al. (2025) construct datasets using LLM-generated or expert-written natural language requirements, failing to capture the diversity and authenticity of real-world user expressions. Second, reference-based benchmarks Beltramelli (2018); Vu et al. (2025); Li et al. (2024); Gui et al. (2024); Yun et al. (2024); Gui et al. (2025) require LLMs to reproduce web apps from provided screenshots or sketches, thus lacking the generality needed to evaluate open-ended user requirements. Third, arena-style benchmarks LMArena.ai (2025); Xiao et al. (2025) collect real-world user requirements and rank LLMs through large-scale human voting in blind evaluations Chiang et al. (2024). While these results align with human preferences, such benchmarks depend heavily on manual annotations and fail to deliver fine-grained, quantitative evaluations for a deep analysis.
@@ 2.1 LLMs for Web App Generation
The LLMs for web app generation mainly fall into three categories. The first category is foundation models, including general-purpose models (e.g., GPT5 OpenAI (2025), Claude Anthropic (2025b), Gemini Team et al. (2023), DeepSeek Liu et al. (2024), Qwen Yang et al. (2025), Doubao ByteDance (2025b)) and code-specialized models (e.g., Deepseek-Coder Guo et al. (2024), Qwen-Coder Hui et al. (2024), StarCoder Li et al. (2023)). These models process multi-language code synthesis, long-context reasoning, and iterative repair via execution feedback. The second category is IDE/CLI-centric coding agents (e.g, Cursor Anysphere (2025), Trae ByteDance (2025c), Windsurf Windsurf (2025), and Claude Code Anthropic (2025a)). These agents support manipulating local projects through project bootstrapping, dependency resolution, multi-file editing, and command-line tool control. The third category is chat-based coding agents (e.g., Manus Manus (2025), MiniMax agent MiniMax (2025a), Doubao Coding ByteDance (2025a), and Lovable Lovable (2025)).
@@ 2.2 Benchmarks for Web App Generation
WebGen-Bench Lu et al. (2025) Text 101 LLM + human experts Yes LLM as a judge + test case execution Test cases
FullFront Sun et al. (2025) Text + Image 1,800 LLM Yes Code & visual similarity + LLM as a judge Code + Images
FrontendBench Zhu et al. (2025) Text 148 LLM Yes Test case execution Test cases
Web-Bench Xu et al. (2025) Text 100 Human experts Yes Test case execution Test cases
WebDev Arena LMArena.ai (2025) Text 10,501 Real user requirements No Pairwise manual labeling -
Design Arena Arena (2025) Text - Real user requirements No Pairwise manual labeling -
ArtifactsBench Zhang et al. (2025) Text 1,825 LLM + human experts Yes LLM as a judge Checklist
@@
Web app generation has long been an active research topic. Beginning with pix2code Beltramelli (2018), a large body of benchmarks and datasets Bhathal and Gupta (2025); Vu et al. (2025); Li et al. (2024); Gui et al. (2024); Yun et al. (2024); Gui et al. (2025) focus on generating web apps from screenshots or sketches. These approaches typically conduct evaluation by comparing the generated web app code or rendered screenshots against the corresponding ground-truth code or original screenshots.
@@ 3.1 Data Collection
Next, we let each requirement be reviewed by three human experts to remove requirements that are incomprehensible, missing supplementary materials (e.g., images), or do not apply to native HTML scenarios. We further conduct text-level and semantic-level de-duplication using MinHash Broder (1997) and MiniLM Wang et al. (2020) semantic embedding, following ArtifactsBench Zhang et al. (2025). Our filtered dataset retains 1,572 requirements.
@@ 3.2 Dataset Statistics
We perform detailed and multi-dimensional classifications of each requirement to enable an in-depth analysis of the characteristics and distribution of our dataset. The classification results are shown in Table 2, with all tags manually labeled by human experts assisted by an LLM (Gemini-2.5-pro).
@@
The artifact complexity dimension spans from highly simple to highly complex outputs, with most samples falling into medium (658) and simple (433) levels, showing balanced difficulty for model evaluation. In terms of input modality, our dataset mainly includes text-only requirements (1,413), complemented by text with images (123) and text with URLs (36), allowing evaluation of both uni-modal and multi-modal understanding abilities. It is worth noting that, unlike the task of generating code from screenshots Beltramelli (2018); Gui et al. (2025), our user requirements can include images and URLs that are intended to serve as page content rather than as reference designs.
@@ 4 Evaluation Metrics
To comprehensively assess the quality of web app artifacts generated by LLMs, we design a set of evaluation metrics from multiple perspectives. Inspired by practices in the text-to-image domain Hartwig et al. (2025), our evaluation considers two major aspects: general quality and alignment quality. These aspects are further divided into nine perspectives encompassing 24 detailed evaluation metrics, jointly providing a comprehensive evaluation. The metrics are defined based on public standards (e.g., Google Lighthouse Google (2025), Lint Hint (2025); Stylelint (2025); ESLint (2025), W3C design principles Consortium (2025)) and internal guidelines from our industrial partner (e.g., visual design and copywriting standards). We list all metrics in Table 3, with detailed information provided in Appendix A. Each metric produces a quantitative score ranging from 0 to 100, where a higher score indicates better quality.
@@ 4.3 Weight Assignment
After defining evaluation metrics, a crucial process is to combine the outcome scores in a meaningful way to generate an overall score, reflecting human preference. Our process of generating the overall score is shown in Figure 3. Existing studies Lin et al. (2025); Zhang et al. (2025) typically assign weights for each metric either uniformly or heuristically. However, real-world users do not treat all metrics equally and can prioritize certain perspectives (e.g., emphasizing code quality while paying less attention to maintainability). Consequently, heuristic weighting neither aligns with user priorities nor ensures fairness.
@@
To obtain preference-aligned weights for each metric, we conduct an internal survey within the company of our industrial partner. Specifically, we ask each participant to rank the 24 metrics according to their perceived importance. To reduce the cognitive burden on participants, each of them is first asked to rank the nine perspectives listed in Table 3, and then to rank the metrics within each perspective. During a three-day survey period, our questionnaire receives 1,076 views and 899 responses, corresponding to a response rate of 83.55%. We further filter the responses by completion time, retaining only those that take more than two minutes to complete. As a result, we obtain 141 valid responses. According to a recent study Memon et al. (2020), a sample size of more than 100 responses can be considered as sufficient for ranking and regression analyses. Among the valid responses, there are 21 data scientists, 19 product managers, 22 legal personnel, 23 front-end/back-end developers, 2 designers, 39 operations staff, and 15 quality assurance personnel, according to their user personas.
We further apply the Borda Count Young (1974) algorithm to extract weights from responses. This algorithm assigns a score to each item (perspective or metric) based on its position in each participant’s ranking. The weight of each item is then obtained by summing its scores across all participants and normalizing by the total score over all items. Each metric’s weight is calculated by multiplying its own weight by the weight of the perspective it belongs to. The weights of all metrics sum to 1. The resulting weights are shown in Figure 4.
The overall score of each model is calculated by summing the z-scores Al Shalabi et al. (2006) of all metrics, averaged over the 1,572 samples in the dataset, and weighted by the corresponding metric weights. We adopt z-scores to ensure that the scores of different metrics are on a comparable scale with consistent discriminability, thereby emphasizing the effect of the weights. The z-scores are calculated by $z_{i,j}=\frac{x_{i,j}-\mu_{j}}{\sigma_{j}}$, where $x_{i,j}$ denotes the raw score of the $i$-th sample on the $j$-th metric, and $\mu_{j}$ and $\sigma_{j}$ denote the mean and standard deviation of this metric over all samples, respectively.
@@ 5.1 Settings
We evaluate 12 representative LLMs and 2 LLM-based agents on WebCoderBench. The selected models span multiple families (e.g., DeepSeek Liu et al. (2024), Qwen Yang et al. (2025), Gemini Team et al. (2023), GLM Zeng et al. (2025), GPT OpenAI (2025)), multiple versions (e.g., base, instruct, thinking, coder), multiple modalities (e.g., text-only, multi-modal), and multiple capability types (e.g., standard Chen et al. (2025); MiniMax (2025b), agentic Manus (2025); MiniMax (2025a)).We invoke each LLM using its standard APIs with the recommended parameters and identical system prompts. For LLM-based agents, we manually enter prompts and requirements through their web interfaces to obtain the generated artifacts, and further constrain their outputs to native HTML by appending specific control prompts. Due to labor constraints, we collect artifacts for 165 requirements from these agents, corresponding to approximately one-tenth of the full dataset. To enable uni-modal LLMs (e.g., DeepSeek) to handle multi-modal requirements, we use Gemini-2.5-pro to generate textual descriptions for each image and provide these descriptions as input to the uni-modal LLMs. We let each model generate a single artifact for each requirement.
@@ 5.2 RQ1: Main Results
Compared to coder models (e.g., Qwen3-Coder-Plus and GPT-5-Codex), generalist models (e.g., Qwen3-Instruct and GPT-5-High) present better effectiveness, showing the task of web app generation requires not only coding ability, but also general understanding of requirements and real-world knowledge. This finding also aligns to ArtifactsBench Zhang et al. (2025).
@@ 6 Conclusion
Finally, metric correctness and stability can influence the evaluation. For rule‑based metrics, we manually inspect 50 scoring instances per metric and conduct additional code reviews. For LLM‑as‑a‑judge metrics, we run each metric three times on 50 requirements and use the Mann–Whitney U test Mann and Whitney (1947) to verify that score variance across runs is substantially lower than variance across models. We follow prior work Sun et al. (2025) for prompt design and use different LLMs for different metrics to mitigate preference leakage and bias Chehbouni et al. (2025); Li et al. (2025); Sheng et al. (2025).
@@ F.5 Prompts for Evaluation Metrics
Prompts for the evaluation metric of Content Alignment: **Role-play:** You are a senior requirements analyst, project manager, and a user of web front-end applications. You have just received a web front-end application developed based on given requirements. It is still in the early stages of development. Your task is to determine whether the developer has correctly understood the user’s requirements and to judge at a high level whether the given web page code meets the user-specified requirement points, without worrying about the correctness of the specific implementation. **Task Goal:** Analyze the web front-end application code and determine if it meets the given requirement points. **Input:** **Web page code** and **requirement points** information **Evaluation Criteria:** Please analyze from the user’s perspective whether the code implements the specified requirement points based on the following dimensions: **Content Semantics:** Does the content on the page conform to the semantics specified in the user requirements? **Specified Data:** Does the content on the page include the specific data that the user specified to include? (e.g., image links, data source links, text copy, video/audio links, etc.) **Output:** Please output your judgment and analysis in JSON format, with the following structure: [{ "content_requirement": "<Requirement Point>", "code_snippet": "<Web Front-end Application Code> (The maximum output length is 1000 characters.)", "is_implemented": <true/false>, "implementation_analysis": "<A concise analysis of how the code implements the requirement point>", "confidence_score": <A confidence score from 0-1, indicating your level of certainty, with 1 being the highest> }] **Example:** **Input:** 1. **Code Snippet:** <div>Welcome</div> <img src="2601.02430v2/logo.png"> <a href="/home"></a> 2. **Requirement Points:** ["Display a welcome text.", "Display a company logo, using logo.png.", "Provide a link to return to the homepage."] **Expected Output (JSON format):** Note: The final output should only contain the JSON content, do not wrap it in Markdown’s JSON format. [{ "content_requirement": "Display a welcome text.", "code_snippet": "<div>Welcome</div>", "is_implemented": true, "implementation_analysis": "The code uses a ‘<div>’tag to contain the text Ẅelcome,̈ thus implementing the requirement to display a welcome text.", "confidence_score": 0.8 }, { "content_requirement": "Display a company logo.", "code_snippet": "<img src=l̈ogo.png>̈", "is_implemented": true, "implementation_analysis": "The code uses an ‘<img>’ tag to attempt to display the user-specified image named l̈ogo.png.̈", "confidence_score": 1.0 }, { "content_requirement": "Provide a link to return to the homepage.", "code_snippet": "<a href=/̈home>̈</a>", "is_implemented": false, "implementation_analysis": "The code uses an ‘<a>’ tag to create a link pointing to /̈home,̈ but there is no text on the link, which fails to indicate to the user that it is a link to return to the homepage.", "confidence_score": 0.9 }]